In a parallelogram, diagonals RT and SW intersect at point A. If SA=x-13 and AW= 2x-37, find SW.
Answer by ikleyn(52810)

.
In a parallelogram, diagonals RT and SW intersect at point A. If SA=x-13 and AW= 2x-37, find SW.
~~~~~~~~~~~~~~~~~~~~

In any parallelogram, diagonals bisect each other at the intersection point.

So, we can write  SA = AW,  or

    x-13 = 2x - 37.


Simplify and find x

    37 - 13 = 2x - x

      24    =   x

       x    =   24.


Then  SA = x-13 = 24 - 13 = 11;  AW = SA = 11;  SW = SA + AW = 11 + 11 = 22.    ANSWER

Solved.